What is the value of x1 in the following equations? X1 - 6x2 + 2x3 = 6 - 2x2 + 5x3 = 31 4x1 - x2 + x3 = 21 O2 04 O 7 O1
Answer #1

Soli Given: 6x2 + 203 6 -(i) - 2x2 +5x3 = 31 (ii) ux - X2 + x3 liii) Multiply equation (i) by 4 then subtract equation (int)Thus My 3+ 23x2 7 7 ice h = 7 =7 NOW the value of and da in equation (i) to get the substitute х, value of a 6x₂ + 202 = 6 -